The area of a circular rug is about 452.2 ft2. What is the approximate radius of the rug?
Found 2 solutions by nyc_function, checkley77:
Answer by nyc_function(2741)

Area of circle = pi times (radius)^2
452.2 = (3.14)(r)^2
(452.2)/3.14 = r^2
144.0127389 = r^2
We now take the square root of both sides.
12.00053077 = radius
This means the radius is about 12 feet.

Answer by checkley77(12844)

Area=pir^2
452.2=3.14r^2
r^2=452.2/3.14
r^2=144
r=sqrt144
r=12 ft. ans. for the radius.